The loaded cab of an elevator has a mass of 3.0 103 kg and moves 240 m up the shaft in 23 s at constant speed. At what average rate does the force from the cable do work on the cab?
Answer #1

Given that the mass of elevator cab is m = 3.0*10^3 kg
The distance moved by cab is h = 240 m
Time taken is t= 23s
-----------------------------------------------------------------
From work energy theorem
Work done by the gravitational force is
W = mgh
Then the power is
P = work done /time
P = mgh / t
P = (3.0*10^3 kg)(9.8 m/s^2)(240 m) / (23 s)
P = 306.78*10^3 watt